Ellen wishes to mix candy worth $2.15 per pound with candy worth $3.95 per pound to form 20 pounds of mixture worth $3.50 per pound. How many pounds of the more expensive candy should she use?
Answer by checkley77(12569)

3.95x+2.15(20-x)=3.50*20
3.95x+43-2.15x=70
1.8x=70-43
1.8x=27
x=27/1.8
x=15 pounds of the $3.95 candy is used.
20-15=5 pounds of $2.15 candy is used.
Proof:
3.95*15+2.15*5=3.50*20
59.25+10.75=70
70=70